# Resource-Efficient Federated Multimodal Learning via Layer-wise and Progressive Training

### v1

@@ IV-B Prog-FedMML
V Experiment
V-A Experimental Setup
Dataset. We conduct our experiments for multimodal learning in two settings: an instance discrimination-based setting and a supervised setting with labeled data. We use the COCO [35] dataset for the instance discrimination-based setting. This dataset contains 118,287 images for training and 5,000 images for testing, with each image paired with at least five captions. We evaluate the performance on text-to-image and image-to-text retrieval tasks. For the supervised approach, we use the ADVANCE [9] dataset for audiovisual aerial scene recognition. This dataset consists of 5,075 aerial images accompanied by their respective audio recordings, categorized into 13 different scene classes. The underlying premise is that geographic locations are associated with certain sound events, and utilizing both aerial images and sound can lead to more accurate predictions. We allocate 20% of the data for testing.
We uniformly split the data among the clients for the COCO dataset. We set the batch size to 256 and use $\tau=0.05$ in Eq. 3. For LW-FedMML and Prog-FedMML, we use the learning rates shown in Table III for the image and text encoders at each stage. The learning rate for the projection heads of both modalities is set to 3e-4. For the conventional end-to-end FedMML, the learning rates for the image and text encoders are set to 1e-7 and 1e-5, respectively, while the projection heads have a learning rate of 1e-4. Unless otherwise stated, we use the COCO dataset for our experiments.
[…]
@@ V-A Experimental Setup
(a) Text-to-Image.
(b) Image-to-Text.
(c) Memory usage.
@@
V-B Instance Discrimination-based Setting
In Table IV, we compare the performance of different approaches and the relative resource requirements for a single client. We measure the maximum memory required during training, while FLOPs and communication costs are measured in total across all local training epochs and FL rounds. For FLOPs calculation, we follow the same strategy as in [32], accounting for only one input sample rather than the entire batch. As the number of samples increases, the FLOPs would increase drastically. Table IV demonstrates that LW-FedMML achieves comparable performance to end-to-end FedMML, with a minor drop of $\approx 1\%$ in average recall for text-to-image and image-to-text retrieval tasks, while significantly reducing resource requirements. Specifically, LW-FedMML shows a 2.7×2.7$\times$2.7 × reduction in memory usage, a 2.4×2.4$\times$2.4 × decrease in FLOPs requirements, and 2.3×2.3$\times$2.3 × lower communication costs. Prog-FedMML achieves the best performance, surpassing end-to-end FedMML, which aligns with findings from prior studies [31, 32]. However, the resource requirements of Prog-FedMML are significantly higher than those of LW-FedMML, closely matching those of end-to-end FedMML in terms of memory and communication costs. This could present similar challenges for Prog-FedMML as those faced by end-to-end FedMML, where many clients may fail to meet the resource requirements for participation, thereby preventing their data from being used in model training. Fig. 6 plots the memory usage, FLOPs, and communication costs during the training process for a client. Across all metrics, LW-FedMML consistently shows significantly lower resource requirements compared to FedMML. In contrast, resource demands for Prog-FedMML gradually increase throughout the training, eventually matching those of FedMML.
(a) FedMML
(b) Prog-FedMML
Figure 8: Relative gradient values of linear layers corresponding to the attention output in DistilBERT. (a) In FedMML, the gradient value diminishes as it backpropagates to earlier layers. (b) Conversely, Prog-FedMML maintains improved gradient flow, allowing for more effective updates.
TABLE VI: Comparison between FedMML (end-to-end), Prog-FedMML (progressive), and LW-FedMML (layer-wise) for audiovisual aerial scene recognition on the ADVANCE dataset. We also compare the relative resource requirements of each approach for a client.
@@ V-B Instance Discrimination-based Setting
Memory FLOPS Comm. Precision Recall F1
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 99.82 94.49 96.93
Prog-FedMML 1.00$\times$ 0.91$\times$ 0.95$\times$ 100.00 95.76 97.75
LW-FedMML 0.51$\times$ 0.41$\times$ 0.54$\times$ 99.90 95.17 97.36
@@
Figure 9: Computational and communication resources required for a client on the ADVANCE dataset. (a) Memory requirement. (b) FLOPs consumption. (c) Communication cost.
V-C Tracking Gradients
We investigate the potential cause behind why progressive training outperforms conventional end-to-end training. Using the text encoder (i.e., DistilBERT) with six transformer blocks, we examine the gradient values of linear layers corresponding to the attention output in each transformer block. For Prog-FedMML, we measure the gradient value for each layer at its corresponding stage. The gradient values relative to the layer in the last transformer block are illustrated in Fig. 8. For FedMML, the gradient value diminishes as it backpropagates to earlier layers. This diminishing pattern is not observed in Prog-FedMML. Although not yet thoroughly investigated, increasing the model depth step-by-step may enable earlier layers to receive more effective gradient updates by mitigating small updates caused by vanishing gradients or large divergences due to exploding gradients. We believe this is a key factor contributing to the superior performance of progressive training compared to conventional end-to-end training.
V-D Eligible Clients
@@
TABLE VII: Performance comparison across varying levels of data heterogeneity on the ADVANCE dataset.
$\beta=50$ $\beta=5$ $\beta=0.5$ $\beta=0.05$
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
FedMML 99.92 95.07 97.32 99.92 95.07 97.32 99.82 94.49 96.93 96.70 74.17 81.76
Prog-FedMML 100.00 96.49 98.13 100.00 96.43 98.11 100.00 95.76 97.75 95.31 76.17 82.80
LW-FedMML 100.00 95.42 97.54 99.92 95.41 97.49 99.90 95.17 97.36 96.60 74.51 81.36
TABLE VIII: Performance comparison on the ADVANCE dataset with varying numbers of participating clients $N_{\text{p}}$ at each FL round. The total number of clients $N$ is set to 40, and the total number of FL communication rounds $R$ is set to 120.
$N_{\text{p}}/N$ $30/40$ $(75\%)$ $20/40$ $(50\%)$ $10/40$ $(25\%)$
Precision Recall F1 Precision Recall F1 Precision Recall F1
FedMML 99.87 93.27 96.29 99.87 93.47 96.39 99.87 93.71 96.50
Prog-FedMML 99.95 92.93 96.08 99.90 92.84 95.98 99.95 93.27 96.22
LW-FedMML 99.90 90.44 94.59 99.90 90.54 94.62 99.90 90.54 94.62
V-E Batch Size
Even though a smaller batch size can reduce memory usage, instance discrimination-based approaches benefit from larger batch sizes due to the increased number of samples. We compare the performance of different training approaches with various batch sizes in Figs. 7a and 7b, while Fig. 7c illustrates the maximum memory usage for each setting. The results indicate that smaller batch sizes lead to reduced performance across all methods. Fig. 7c shows a sharp increase in memory requirements for end-to-end FedMML and Prog-FedMML even as the batch size grows from 32 to 256. In contrast, the memory requirements for LW-FedMML remain relatively low across the same batch sizes. Fig. 7 also demonstrates that LW-FedMML can achieve the same performance as end-to-end FedMML when using a larger batch size (i.e., 1024) while still using less memory than FedMML with a batch size of 256. Moreover, increasing the batch size in LW-FedMML only impacts the memory usage, allowing it to maintain its advantage of lower communication costs.
V-F Supervised Setting
For the ADVANCE dataset, we allocate FL rounds to later stages for Prog-FedMML, as we do for LW-FedMML in Table II. Table VI compares the performance of different approaches and relative resource requirements for a single client. The measurement of resource requirements follows the same approach as in the instance discrimination-based setting. The table shows that Prog-FedMML achieves the best performance, while LW-FedMML maintains competitive performance across various metrics compared to FedMML in a supervised multimodal learning setting. In Fig. 9, we can observe that LW-FedMML significantly lowers the resource requirements compared to FedMML, while Prog-FedMML exhibits similar resource demands as FedMML.
V-G Data Heterogeneity
@@
Data heterogeneity is one of the main challenges in FL [20]. Variations in device usage patterns, geographical locations, and other factors influence the data distribution among clients. To control the strength of data heterogeneity within the clients’ local datasets, we adjust the $\beta$ parameter in the Dirichlet distribution. A smaller $\beta$ value corresponds to a higher level of data heterogeneity, and vice versa. Fig. 10 illustrates the data distribution among clients with different $\beta$ values for the ADVANCE dataset, and Table VII shows their results. These results indicate that the performance of all approaches declines as the level of data heterogeneity increases. Nonetheless, in each setting, Prog-FedMML achieves the best performance for most metrics, while LW-FedMML consistently maintains good performance comparable to the end-to-end FedMML.
V-H Client Participation Ratio
@@
In an FL environment with $N$ total clients, it is often impractical for all $N$ clients to simultaneously participate in every FL training round. Therefore, at each FL round, we randomly select a subset of clients to participate in the training. For this experiment, we fixed the total number of clients to $N=40$ and set the number of participating clients $N_{\text{p}}$ to be 10, 20, and 30, while using the ADVANCE dataset. The total number of FL communication rounds $R$ is set to 120. For both LW-FedMML and Prog-FedMML, we allocate the FL rounds for each stage $R_{s}$ as $[0,0,0,20,30,70]$. The results in Table VIII demonstrate that LW-FedMML can sustain reasonable performance even when the client participation ratio is low. Prog-FedMML achieves slightly higher precision but shows a minor drop in recall and F1 metrics compared to FedMML.
@@ VI Conclusion
Resource constraints are a significant challenge in the FL process, mainly because FL clients are often edge devices with limited capabilities. This challenge is exacerbated in FL scenarios involving multimodal data due to the complexity of the models required for processing such data. Many clients cannot meet the resource requirements for conventional end-to-end training approaches in these scenarios, resulting in their valuable data being excluded from the training process. To address this issue, we introduce LW-FedMML, a layer-wise training approach designed to significantly reduce the resource burden on clients. Our experiments demonstrate that LW-FedMML can compete with conventional end-to-end FedMML while achieving up to 2.7×2.7$\times$2.7 × lower memory usage, 2.4×2.4$\times$2.4 × FLOPs requirements, and 2.3×2.3$\times$2.3 × cheaper communication costs on the COCO dataset. These findings highlight the potential of LW-FedMML to enable efficient and scalable multimodal learning in resource-constrained FL environments. Additionally, we explore Prog-FedMML, a progressive training approach for federated multimodal learning. While Prog-FedMML has higher resource requirements than LW-FedMML, it can outperform end-to-end FedMML in most cases.

### v2

@@ V-B Instance Discrimination-based Setting
(a) FedMML
(b) Prog-FedMML
Figure 7: Relative gradient values of linear layers corresponding to the attention output in DistilBERT. (a) For FedMML, the gradient value diminishes as it backpropagates to earlier layers. (b) Conversely, Prog-FedMML maintains improved gradient flow, allowing for more effective updates.
V-C Tracking Gradients
We investigate the underlying factors that might contribute to the superior performance of progressive training compared to conventional end-to-end training. Using the text encoder (i.e., DistilBERT) with six transformer blocks, we examine the gradient values of linear layers corresponding to the attention output in each transformer block. We use the same batch of samples for measuring the gradient values. For FedMML, we measure the gradients before the training when errors and gradient updates are expected to be the largest. For Prog-FedMML, we measure the gradient value of each layer at the first round of its corresponding stage. The gradient values relative to the layer in the last transformer block are illustrated in Fig. 7. For FedMML, the gradient value diminishes as it backpropagates to earlier layers. This diminishing pattern is not observed in Prog-FedMML. Although not yet thoroughly investigated, increasing the model depth step-by-step may enable earlier layers to receive more effective gradient updates by mitigating small updates caused by vanishing gradients or large divergences due to exploding gradients. We believe this is a key factor contributing to the superior performance of progressive training compared to conventional end-to-end training.
V-D Eligible Clients
@@
(a) Memory usage.
(b) Text-to-Image.
(c) Image-to-Text.
@@
TABLE VI: Time and energy consumption for a client.
Number of Samples 1000 10,000 100,000
Time (s) Energy (Wh) Time (s) Energy (Wh) Time (s) Energy (Wh)
FedMML 326 23 2080 144 19865 1380
Prog-FedMML 312 21 1963 136 18490 1284
LW-FedMML 249 17 1107 77 9487 659
V-E Batch Size
Even though a smaller batch size can reduce memory usage, instance discrimination-based approaches benefit from larger batch sizes due to the increased number of samples. We compare the maximum memory usage of each training approach with different batch sizes in Fig. 8a, while Figs. 8b and 8c compare their performance. The results indicate that smaller batch sizes lead to reduced performance across all methods. Fig. 8a shows a sharp increase in memory requirements for end-to-end FedMML and Prog-FedMML even as the batch size grows from 32 to 256. In contrast, the memory requirements for LW-FedMML remain relatively low across the same batch sizes. Fig. 8 also indicates that LW-FedMML with a batch size of 1024 uses less memory than end-to-end FedMML with a batch size of 256, while achieving the same level of performance. Moreover, increasing the batch size in LW-FedMML only impacts the memory usage, allowing it to maintain its advantage of lower communication costs.
V-F Time and Energy Consumption
[…]
TABLE VII: Comparison between FedMML (end-to-end), Prog-FedMML (progressive), and LW-FedMML (layer-wise) for audiovisual aerial scene recognition on the ADVANCE dataset. We also compare the relative resource requirements of each approach for a client.
@@ V-F Time and Energy Consumption
Memory FLOPs Comm. Precision Recall F1 Accuracy
Concat
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 75.91 73.79 74.78 95.46
Prog-FedMML 1.00$\times$ 0.91$\times$ 0.95$\times$ 74.79 72.91 73.80 95.76
LW-FedMML 0.63$\times$ 0.41$\times$ 0.54$\times$ 74.19 72.17 73.11 95.17
Concat + Add
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 73.18 71.30 72.16 94.58
Prog-FedMML 1.00$\times$ 0.91$\times$ 0.95$\times$ 81.12 79.64 80.33 96.68
LW-FedMML 0.61$\times$ 0.41$\times$ 0.54$\times$ 74.48 72.53 73.44 95.02
Concat + Self-Attention [43]
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 71.85 69.70 70.71 94.98
Prog-FedMML 1.00$\times$ 0.91$\times$ 0.95$\times$ 76.35 74.45 75.34 95.56
LW-FedMML 0.60$\times$ 0.41$\times$ 0.54$\times$ 74.01 72.00 72.93 94.82
Concat + Co-Attention [24]
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 69.79 67.32 68.48 94.39
Prog-FedMML 1.00$\times$ 0.91$\times$ 0.95$\times$ 72.03 69.76 70.80 94.63
LW-FedMML 0.61$\times$ 0.41$\times$ 0.54$\times$ 73.62 71.41 72.40 94.43
@@
Figure 9: Computational and communication resources required for a client when using the ADVANCE dataset with concatenation as the fusion technique. (a) Memory requirement. (b) FLOPs consumption. (c) Communication cost.
V-G Supervised Setting
For the ADVANCE dataset, we allocate more FL rounds to later stages of Prog-FedMML, as we do for LW-FedMML in Table II. We use various fusion techniques, including concatenation, addition, self-attention [43], and co-attention [24] to combine the image and audio projections for the prediction. Table VII compares the performance of different approaches and relative resource requirements for a single client. We follow the same approach as in the instance discrimination-based setting for measuring resources. The results show that Prog-FedMML achieves the best performance in terms of accuracy, while LW-FedMML either maintains competitive performance or outperforms FedMML across various metrics. There is no significant variation in terms of relative resource requirements when using different fusion techniques. Fig. 9 plots the resource requirements of a client when using concatenation as the fusion approach. The plots indicate that LW-FedMML significantly lowers the resource requirements compared to FedMML, while Prog-FedMML exhibits similar resource demands as FedMML.
@@ V-G Supervised Setting
V-H Data Heterogeneity
@@
Data heterogeneity is one of the main challenges in FL [20]. Variations in device usage patterns, geographical locations, and other factors influence the data distribution among clients. To control the degree of data heterogeneity within the clients’ local datasets, we adjust the $\beta$ parameter in the Dirichlet distribution. A smaller $\beta$ value corresponds to a higher level of data heterogeneity, and vice versa. Fig. 10 illustrates the data distribution among clients with different $\beta$ values for the ADVANCE dataset, and Table VIII shows their results. These results indicate that the performance of all approaches declines as the level of data heterogeneity increases. Nonetheless, both Prog-FedMML and LW-FedMML consistently maintain strong performance across all levels of data heterogeneity.
TABLE VIII: Performance comparison across varying levels of data heterogeneity on the ADVANCE dataset.
Precision Recall F1 Accuracy
$\beta=50$
FedMML 78.80 77.04 77.84 95.91
Prog-FedMML 80.78 79.21 79.95 96.49
LW-FedMML 77.60 75.63 76.55 95.52
$\beta=5$
FedMML 75.39 73.41 74.33 95.52
Prog-FedMML 80.60 79.10 79.81 96.64
LW-FedMML 77.24 75.39 76.24 95.42
$\beta=0.5$
FedMML 75.91 73.79 74.78 95.46
Prog-FedMML 74.79 72.91 73.80 95.76
LW-FedMML 74.19 72.17 73.11 95.17
$\beta=0.05$
FedMML 52.46 46.28 48.37 73.98
Prog-FedMML 52.33 46.44 48.25 74.61
LW-FedMML 54.96 48.68 50.74 74.12
V-I Client Participation Ratio
@@
In an FL environment with $N$ total clients, it is often impractical for all $N$ clients to simultaneously participate in every FL training round. Therefore, at each FL round, we randomly select a subset of clients to participate in the training process. Using the ADVANCE dataset, we fixed the total number of clients to $N=50$ and set the number of participating clients $N_{\text{p}}$ to be 35, 25, and 15. The total number of FL communication rounds $R$ is set to 300. For both LW-FedMML and Prog-FedMML, we allocate 50 FL rounds for each stage $s$. The learning rate is set to 1e-4 for all modules in FedMML and Prog-FedMML. For LW-FedMML, we set the learning rate of image and audio encoders to 1e-4, and to 1e-3 for the remaining modules. The results in Table IX demonstrate that LW-FedMML can maintain decent performance even when the client participation ratio is low, despite having significantly lower resource requirements compared to FedMML. Meanwhile, Prog-FedMML achieves the best performance in most cases.
We also consider an experiment where each client is assigned a dropout probability, simulating instances where clients may leave the training process due to communication failures, bandwidth limitations or other operational constraints. In this setting, we set $N=50$ and examine two dropout probabilities: 25% and 50%. The results for this setting are shown in Table X.
TABLE IX: Performance comparison on the ADVANCE dataset with different numbers of participating clients $N_{\text{p}}$ at each FL round. The total number of clients $N$ is fixed at 50, while $N_{\text{p}}$ is set to 35, 25, and 15.
Precision Recall F1 Accuracy
$N_{\text{p}}/N=35/50$ $(70\%)$
FedMML 72.89 70.42 71.54 94.29
Prog-FedMML 77.21 74.78 75.81 94.83
LW-FedMML 70.52 68.38 69.36 94.00
$N_{\text{p}}/N=25/50$ $(50\%)$
FedMML 72.50 70.12 71.20 94.19
Prog-FedMML 76.43 73.76 74.90 94.11
LW-FedMML 72.40 70.16 71.18 95.27
$N_{\text{p}}/N=15/50$ $(25\%)$
FedMML 73.80 71.50 72.55 94.68
Prog-FedMML 76.84 73.86 75.06 94.30
LW-FedMML 70.70 68.15 69.30 93.90
TABLE X: Performance comparison on the ADVANCE dataset with varying client dropout rates.
Precision Recall F1 Accuracy
dropout rate = 25%
FedMML 72.37 69.85 70.99 94.29
Prog-FedMML 77.08 74.70 75.72 94.79
LW-FedMML 72.53 70.35 71.33 95.07
dropout rate = 50%
FedMML 75.10 72.83 73.86 94.68
Prog-FedMML 78.52 76.23 77.22 95.02
LW-FedMML 72.60 70.38 71.36 94.04
TABLE XI: ResNet-18 as image encoder.
Resource Requirements
Memory FLOPs Comm. Precision Recall F1 Accuracy
FedMML 1.00$\times$ 1.00$\times$ 1.00$\times$ 69.97 67.69 68.74 94.39
Prog-FedMML 1.00$\times$ 0.80$\times$ 0.85$\times$ 69.38 66.72 67.88 93.65
LW-FedMML 0.66$\times$ 0.41$\times$ 0.57$\times$ 69.56 67.13 68.21 93.55
TABLE XII: Training from scratch.
Resource Requirements
Rounds $R$ Memory FLOPs Comm. Precision Recall F1 Accuracy
Same Number of Rounds
FedMML 300 1.00$\times$ 1.00$\times$ 1.00$\times$ 46.45 39.47 41.91 76.69
Prog-FedMML 300 1.00$\times$ 0.59$\times$ 0.78$\times$ 47.18 39.90 42.27 74.83
LW-FedMML 300 0.71$\times$ 0.31$\times$ 0.54$\times$ 41.02 33.32 36.23 75.81
Same Comm. Cost
FedMML 161 1.00$\times$ 1.00$\times$ 1.00$\times$ 44.90 35.92 38.53 68.73
Prog-FedMML 228 0.92$\times$ 0.66$\times$ 1.00$\times$ 49.52 41.11 43.66 70.59
LW-FedMML 300 0.71$\times$ 0.57$\times$ 1.00$\times$ 41.02 33.32 36.23 75.81
V-J ResNet as Image Encoder
When using model architectures other than transformers, appending new layers to the encoder at each stage can alter the encoder output dimensions. To investigate this, we apply ResNet-18 [44] as the image encoder and DistilAST as the audio encoder on the ADVANCE dataset, setting the number of stages to $S=4$. As an additional ResNet block is introduced to the image encoder at each stage, the dimension of the output image representation changes. Specifically, the output dimensions of the image encoder across the four stages are (64, 128, 256, 512), while the input dimension to the projection head remains fixed at 512. Therefore, we introduce a linear layer to reshape the encoder output whenever there is a dimensionality mismatch. This linear layer is discarded after its corresponding stage. For DistilAST, we introduce two transformer blocks at each of the first two stages, followed by one transformer block at each of the last two stages, resulting in a total of six transformer blocks.
The results are shown in Table XI. For all approaches, the learning rates are set as follows: 1e-7 for the image encoder, 1e-5 for the audio encoder, and 1e-4 for all other modules. The total number of FL communication rounds is 90, with rounds for each stage allocated as (10, 20, 25, 35) for both LW-FedMML and Prog-FedMML. Both Prog-FedMML and LW-FedMML demonstrate comparable performance to end-to-end FedMML, with LW-FedMML offering the additional advantage of significantly lower resource consumption.
V-K Training from Scratch
Using a similar setup to Section V-I, we conduct the training without pre-trained weights. Specifically, we use the setting where the number of participating clients in each round is set to $N_{\text{p}}=15$. We also keep other settings except that for LW-FedMML, the learning rate for the image and audio encoders is set to 1e-3, and to 1e-4 for other modules. Table XII presents the results under two scenarios: one where the number of FL communication rounds is the same for all approaches, and another where the communication cost is the same. Under the same communication cost, LW-FedMML achieves higher accuracy than FedMML while also using less memory and fewer FLOPs. The results in Table XII indicate that both Prog-FedMML and LW-FedMML can also operate without using the pre-trained weights.
@@ VI Conclusion
Resource constraints are a significant challenge in the FL environment, mainly because FL clients are often edge devices with limited capabilities. This challenge is exacerbated in FL scenarios involving multimodal data due to the complexity of the models required for processing such data. Many clients cannot meet the resource requirements for conventional end-to-end training approaches in these scenarios, resulting in their valuable data being excluded from the training process. To address this issue, we introduce LW-FedMML, a layer-wise training approach designed to significantly reduce the resource burden on clients. Our experiments demonstrate that LW-FedMML can compete with conventional end-to-end FedMML while achieving up to 2.7×2.7$\times$2.7 × lower memory usage, 2.4×2.4$\times$2.4 × FLOPs requirements, and 2.3×2.3$\times$2.3 × cheaper communication costs on the COCO dataset. These findings highlight the potential of LW-FedMML to enable efficient and scalable multimodal learning in resource-constrained FL environments. Additionally, we explore Prog-FedMML, a progressive training approach for federated multimodal learning. While Prog-FedMML has higher resource requirements than LW-FedMML, it has the potential to outperform end-to-end FedMML in most cases.
Appendix A Supervised LW-FedMML
Algorithms 3 and 4 describe the details for the supervised setting of LW-FedMML. In the supervised setting, the local dataset $D^{n}$ on each client $n$ contains ground truth labels $y$. The prediction head $H_{\text{sup}}$ generates predictions $\hat{y}$ and the cross-entropy loss is calculated between $y$ and $\hat{y}$. In addition to the active layers and the projection heads, the clients also exchange $H_{\text{sup}}$ with the server in the supervised setting.
Figure 11: Comparison between: (a) FedMML (end-to-end), (b) Prog-FedMML (progressive), and (c) LW-FedMML (layer-wise) for the instance discrimination-based setting.
Algorithm 3 Supervised LW-FedMML (Server-side)
Input: encoder $F_{a}^{0}$ and projection head $H_{a}$ for modality $a$, encoder $F_{b}^{0}$ and projection head $H_{b}$ for modality $b$, prediction head $H_{\text{sup}}$, number of stages $S$, number of rounds $R^{s}$ where $s\in[1,S]$
Output: $F_{a}^{S}$, $F_{b}^{S}$, $H_{a}$, $H_{b}$, $H_{\text{sup}}$
1:Server executes:
2:for stage $s=1,2,\dots,S$ do
3:     Initialize new layers: $L_{a}^{s}$ and $L_{b}^{s}$
4:     for round $r=1,2,\dots,R^{s}$ do
5:         for client $n=1,2,\dots,N$ in parallel do
6:              $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n},H^{n}_{\text{sup}}\leftarrow$
7:                               $\text{Train}(L_{a}^{s},L_{b}^{s},H_{a},H_{b},H_{\text{sup}})$
8:              $w^{n}=\frac{|D^{n}|}{|D|}$, where $D=\bigcup_{n=1}^{N}D^{n}$
9:         end for
10:         $L_{a}^{s}\leftarrow\sum_{n=1}^{N}w^{n}L_{a}^{s,n},\quad H_{a}\leftarrow\sum_{n=1}^{N}w^{n}H_{a}^{n}$
11:         $L_{b}^{s}\leftarrow\sum_{n=1}^{N}w^{n}L_{b}^{s,n},\quad H_{b}\leftarrow\sum_{n=1}^{N}w^{n}H_{b}^{n}$
12:         $H_{\text{sup}}\leftarrow\sum_{n=1}^{N}w^{n}H^{n}_{\text{sup}}$
13:     end for
14:     Distribute $L_{a}^{s}$ and $L_{b}^{s}$ to clients
15:     $F_{a}^{s}\leftarrow$ Attach $L_{a}^{s}$ to $F_{a}^{(s-1)}$
16:     $F_{b}^{s}\leftarrow$ Attach $L_{b}^{s}$ to $F_{b}^{(s-1)}$
17:end for
18:return $F_{a}^{S}$, $F_{b}^{S}$, $H_{a}$, $H_{b}$, $H_{\text{sup}}$
Algorithm 4 Supervised LW-FedMML (Client-side)
Output: $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n}$, $H^{n}_{\text{sup}}$
1:Client executes: $\text{Train}(L_{a}^{s},L_{b}^{s},H_{a},H_{b},H_{\text{sup}})$:
2:$L_{a}^{s,n}\leftarrow L_{a}^{s},\quad H_{a}^{n}\leftarrow H_{a}$
3:$L_{b}^{s,n}\leftarrow L_{b}^{s},\quad H_{b}^{n}\leftarrow H_{b}$
4:$H^{n}_{\text{sup}}\leftarrow H_{\text{sup}}$
5:$F_{a}^{s,n}\leftarrow$ Freeze $[L_{a}^{1},\dots,L_{a}^{(s-1)}]$ and attach $L_{a}^{s,n}$
6:$F_{b}^{s,n}\leftarrow$ Freeze $[L_{b}^{1},\dots,L_{b}^{(s-1)}]$ and attach $L_{b}^{s,n}$
7:for epoch $e=1,2,\dots,E$ do
8:     for each batch $(x_{a},x_{b},y)\in D^{n}$ do
9:         $z_{a}\leftarrow H_{a}^{n}(F_{a}^{s,n}(x_{a}))$
10:         $z_{b}\leftarrow H_{b}^{n}(F_{b}^{s,n}(x_{b}))$
11:         $\hat{z}\leftarrow$ Fuse $(z_{a},z_{b})$
12:         $\hat{y}\leftarrow H^{n}_{\text{sup}}(\hat{z})$
13:         $\mathcal{L}\leftarrow\ell_{\text{CE}}(y,\hat{y})$
14:         $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n},H^{n}_{\text{sup}}\leftarrow$ Update with $\nabla\mathcal{L}$
15:     end for
16:end for
17:Upload $L_{a}^{s,n},L_{b}^{s,n},H_{a}^{n},H_{b}^{n},H^{n}_{\text{sup}}$
Appendix B Comparison with Hierarchical Gradient Blending
We perform a comparison with FedHGB [25], and the results are presented in Table XIII. FedHGB requires labeled data to calculate its blending coefficients. Therefore, we use the ADVANCE dataset under the same settings as in Section V-G, with concatenation as the fusion approach. The resource requirement of FedHGB is similar to that of FedMML.
TABLE XIII: Comparison with FedHGB [25].
Precision Recall F1 Accuracy
FedHGB 60.01 57.32 58.31 81.64
Prog-FedMML 74.79 72.91 73.80 95.76
LW-FedMML 74.19 72.17 73.11 95.17
Appendix C Illustrative Comparison of Different Approaches
In Fig. 11, we visually highlight the key differences between the three approaches: FedMML (end-to-end), Prog-FedMML (progressive), and LW-FedMML (layer-wise).
Appendix D Discussion
D-A Choosing between LW-FedMML and Prog-FedMML
Here, we discuss the decision-making process for selecting between LW-FedMML and Prog-FedMML, as these approaches can significantly influence resource consumption and performance outcomes. The choice between them can be based on the potential client participation rate, which is further influenced by the resource requirements of each method. While measuring exact resource requirements for each approach can be challenging, it is often feasible to estimate them before initiating the training process. This estimation can be as simple as running both training processes on a sample device. If device statistics (e.g., type of device) from the targeted clients (e.g., those located in a certain region) are available, the potential client participation ratio can be estimated based on the resource requirements. Device statistics can be less sensitive than actual training data and are often accessible in practice. This allows for determining the participation gap between LW-FedMML and Prog-FedMML. If the gap is large, meaning many clients cannot support Prog-FedMML, LW-FedMML would be the preferable choice. Conversely, if the gap is small, indicating most clients can handle Prog-FedMML, it would be the more suitable option.
D-B Handling New Modalities
In some cases, a new modality may be introduced to the system after the training stages have been completed. For instance, a sensor with a new modality $c$ could be added to an IoT system that already incorporates modalities $a$ and $b$. For conventional end-to-end training (i.e., FedMML), we can integrate a new encoder $F_{c}$ while freezing the previously trained encoders $F_{a}$ and $F_{b}$, avoiding the need to restart training from scratch. To prevent over-reliance on the pre-trained modalities, regularization techniques such as modality dropout [45] can be applied during the training of $F_{c}$. By dropping out pre-existing modalities $a$ and $b$, the model is forced to rely on new modality $c$, facilitating more effective training of $F_{c}$. This strategy can also be extended to layer-wise and progressive training settings. In LW-FedMML, during each stage $s$, only the layer $L_{c}^{s}$ within the new encoder $F_{c}^{s}$ is trained, while the encoders for the existing modalities (i.e., $F_{a}^{s}$ and $F_{b}^{s}$) are kept frozen. Similarly, in Prog-FedMML, the layers $[L_{c}^{1},\dots,L_{c}^{s}]$ within $F_{c}$ can be trained at each stage $s$, while keeping the other modality branches frozen. In this way, both LW-FedMML and Prog-FedMML can incorporate new modalities without requiring a complete retraining of the entire model.